### v1

# Combine and conquer: model averaging for out-of-distribution forecasting

Abstract
Travel behaviour modellers have an increasingly diverse set of models at their disposal, ranging from traditional econometric structures to models from mathematical psychology and data-driven approaches from machine learning. A key question arises as to how well these different models perform in prediction, especially when considering trips of different characteristics from those used in estimation, i.e. out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models. Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy. To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation. Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation. Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and validation data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
Keywords: choice modelling; forecasting; machine learning; mathematical psychology; mode choice; model averaging
1 Introduction
The travel behaviour modelling literature has focussed extensively on two sorts of models, namely models for inference (henceforth inference models) and models for forecasting (henceforth forecasting models). Inference models aim to understand current travel behaviour (e.g. to recover the value of travel time), while forecasting models aim to predict future travel behaviour in new settings (e.g. due to changes in transport policies, such as toll roads and fuel levies).
Inference models and forecasting models are evaluated differently by analysts. When building an inference model, an analyst is keen to ensure that the model generates behaviourally plausible insights into explanatory factors and their relative impacts. When building a forecasting model, the focus is much more on a model’s ability to generate accurate predictions of behaviour. Importantly, in forecasting, we predict the behaviour in new settings, meaning that the models have to generalise out-of-sample. As some of the settings we wish to make predictions for may be outside the range of level-of-service values that the models have been estimated on, it is important for forecasting models to also generalise out-of-distribution.
Travel behaviour analysts are increasingly interested in a more diverse set of models, covering econometric structures, models from mathematical psychology, and machine learning techniques. With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf. Hagenauer and Helbich, 2017).
Of course, the entire purpose of building a forecasting model is not to replicate choices that have already been made, but to make predictions for new settings. Despite the growing interest in machine learning approaches, the dominant - although not necessarily evidenced-based - view is that forecasting models with a solid behavioural underpinning, such as Random Utility Maximisation (RUM) based discrete choice models, are better equipped to predict behaviour under new settings than are models with a weak or no behaviour underpinning, such as e.g. machine learning models (cf. van Cranenburgh et al., 2022).
However, what is currently less well understood is how to value and incorporate model performance, i.e. the model fit on the empirical data, in the process of developing forecasting models. What is clear is that good model performance is insufficient to establish that a model will make good out-of-distribution predictions. For that, the model must have captured underlying relationships that generalise beyond the observed data. The choice modelling literature has at times recognised that more advanced models that offer a better fit on the empirical data do not necessarily lead to better forecasts (see e.g. Fox et al., 2014). However, the question of how to develop forecasting models considering the model’s performance and behavioural underpinning is still open. On the one hand, the fact that model parameters are estimated by maximising the model performance on the empirical data, i.e. how well they replicate current choices, is somewhat at odds with the aim of forecasting, i.e. to generalise out-of-distribution. On the other hand, it is intuitive that the performance of a model on the empirical data still pertains to relevant information on its ability to generalise out-of-distribution.
The aim of this paper is twofold. First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models. Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance". In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only physical distance.
We employ a neural network for the model averaging. This allows learning potential nonlinear relations between distance and model weights. Using the resulting model averaging structure, we aim not only to develop a flexible tool for combining models but also to gain rule-of-thumb insights for the conditions under which specific sorts of models perform best in terms of out-of-distribution forecasting.
The remainder of this paper is organised as follows. Section 2 presents our proposed model averaging approach. This is followed by the two case studies in Section 3, and conclusions in Section 4.
2 Methodology
As outlined in the introduction, the methodological novelty of the present paper is the use of model averaging for out-of-distribution prediction. In this section, we first discuss the general concept of model averaging before looking into the adaptation required for out-of-distribution prediction.
2.1 Generic model averaging approach
In model averaging (see e.g. Hancock et al., 2020), we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$. Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
Model averaging now combines the insights from these different models into one overall structure. The specification of the model averaging log-likelihood function depends on whether an analyst wishes the weights for individual models to be defined at the person level (i.e. $\pi_{n,m}$ for model $m$ for person $n$) or the observation level (i.e. $\pi_{n,t,m}$ for model $m$ for person $n$ in choice situation $t$). With the former, we would have:
$LL\left(\Theta\right)=\sum_{n=1}^{N}log\sum_{m=1}^{M}\pi_{n,m}L_{n,m}\left(\hat{\Omega}_{m}\right),$ (1)
while, with the latter, we would have:
$LL\left(\Theta\right)=\sum_{n=1}^{N}\sum_{t=1}^{T_{n}}log\sum_{m=1}^{M}\pi_{n,t,m}L_{n,t,m}\left(\hat{\Omega}_{m}\right).$ (2)
with $T_{n}$ choices per person. In this notation, $\Theta$ is a vector of parameters for the model averaging model, while $\hat{\Omega}_{m}$ are the maximum likelihood estimates (MLE) for the vector of parameters $\Omega_{m}$ for model $m$.
The weights are parameterised as being a function of the person ($z_{n}$) and observation ($z_{n,t}$) characteristics, with parameters $\gamma_{m}$ estimated to capture the influence of these characteristics. In particular, we would have:
$\pi_{n,m}\left(\Theta\right)=\frac{e^{\gamma_{m}^{\prime}z_{n}}}{\sum_{l=1}^{M}e^{\gamma_{l}^{\prime}z_{n}}},$ (3)
or
$\pi_{n,t,m}\left(\Theta\right)=\frac{e^{\gamma_{m}^{\prime}z_{n,t}}}{\sum_{l=1}^{M}e^{\gamma_{l}^{\prime}z_{n,t}}},$ (4)
where an appropriate normalisation is needed for $\gamma$, typically setting $\gamma_{m}=0$ for one model $m$.
A reader familiar with latent class models will realise that what is described above is, in essence, a sequential latent class structure. The sequential nature of the process involves first estimating the parameters for the $M$ individual models before then keeping those parameters fixed (at $\hat{\Omega}_{m}$ for model $m$) and estimating weights for each model, as in Equation 1 and 2. In maximising the model averaging likelihood, an analyst thus, in fact, does not require the model parameters for the individual models, only the probabilities of each observed choice.
In applying the model to predict choices out-of-sample, an analyst then again requires the predicted choice probabilities from the individual models first. Let us assume that $P_{i,n,f,m}\left(\hat{\Omega}_{m}\right)$ is the probability of person $n$ choosing alternative $i$ in some forecast scenario $f$, according to model $m$, again calculated at the MLEs. Let us further assume that we have worked with model averaging weights at the observation level, i.e. maximising Equation 2 to obtain MLEs for the model weights at the observation level. We can then use these parameters to calculate the model averaging weights for the forecast scenario, i.e. $\pi_{n,f,m}\left(\hat{\Theta}\right)$. The model averaging probability for choosing alternative $i$ in this new setting would then be given by:
$\hat{P}_{i,n,f,MA}=\sum_{m=1}^{M}\hat{\pi}_{n,f,m}\left(\hat{\Theta}\right)P_{i,n,f,m}\left(\hat{\Omega}_{m}\right).$ (5)
2.2 Model averaging for out-of-distribution prediction
Model averaging typically estimates each model on the entire data and then computes the weights by considering how well each model fits for each of the data points, estimating such weights at the person or observation level. This could, for example, include decision-maker characteristics such as age, gender and income and trip characteristics such as purpose and trip distance.
Let us focus on journey distance in particular. Training a model averaging structure where journey distance, say $d_{n,t}$ for person $n$ and trip $t$, is used as a key characteristic in Equation 4 would allow the model to learn how different sub-models are better suited for predicting trips of specific distances within the bounds of the estimation data. However, if the individual sub-models have been estimated on the entire data, then model averaging by definition cannot gain any insights about how the performance might differ across models out-of-distribution, and thus what weights should be given to them outside the range of distances used in estimation. As a result, the weights will simply be an extrapolation based on the estimates for $\gamma_{n,t,m}$ obtained on the training data. For example, imagine that our model would find that the weight assigned to model $m$ increases with distance. If making predictions out-of-distribution, the model would follow the same trend, and for trips with distances below the lower bound of the estimation data, the weight assigned to model $m$ would decrease, while for trips with distances above the upper bound of the estimation data, the weight assigned to model $m$ would increase.
In the present paper, we propose a different approach that enables us to examine the weight assigned to individual models in out-of-distribution predictions. Again, using distance as an example, let us assume that for the available estimation data, we have minimum and maximum distances $d_{min}$ and $d_{max}$, i.e. $D_{data}=\left[d_{min},d_{max}\right]$. We would then define a narrower interval $D_{train}=\left[d_{a},d_{b}\right]$, where $d_{min}<d_{a}<d_{b}<d_{max}$. In the estimation of the individual sub-models, we then use only a subset of the range of the available data, specifically those observations where $d_{n,t}\in D_{train}$. The model averaging structure on the other hand will be estimated on the entire data, i.e. still maximising e.g. Equation 2. This means that the probabilities for the individual models will be calculated for all trips in the data, but using parameters $\hat{\Omega}_{m}$ obtained from optimisation on only a part of the data (where $d_{n,t}\in D_{train}$). Model averaging learns the role of trip characteristics in determining which model should obtain more weight for a specific trip. These are grouped together in the vector $z_{n,t}$ in Equation 4. By splitting the data in the way described above, we can now define new variables to be used in Equation 4 that help characterise whether a given observation is within $D_{train}$, and if not, how far outside the training data it is located.
The above approach directly allows the model averaging procedure to learn how individual sub-models perform outside the range of distances on which they have been trained and how the weight given to each should, as a result, change the further we move away from that interval. This, in turn, then also means that after estimating the model averaging structure, predictions can be made outside the range $D_{data}$, which is of key interest in travel demand forecasting. The model averaging procedure will have learned how the weight assigned to different models changes outside the estimation data and can then apply that knowledge in out-of-distribution prediction. For example, the model averaging process might determine that models with a stronger behavioural foundation suffer a smaller drop in performance outside the estimation data and will thus gain weight in out-of-distribution prediction.
It should be clear that an analyst will need to make important trade-offs in this process. By increasing the gaps $d_{a}-d_{min}$ and $d_{max}-d_{b}$, the analyst will increase the ability of the model averaging structure to learn about out-of-distribution weights. But this will come at the cost of reducing the width of $D_{train}$ and the size of the sample used for training the individual sub-models, reducing the ability to obtain robust estimates of the influence of level-of-service variables, for example.
Hitherto model-averaging studies have primarily relied on logistic regression models, as shown in (4), often with linear specifications. Because of this, such models are ill-equipped to capture complex relations between explanatory variables used for the model averaging and the likelihood of the model $\pi_{n,m}$. However, when it comes to out-of-distribution prediction using model-averaging, nonlinear effects, e.g., of distance, can be expected. Therefore, in this study, we propose using an MLP for the model averaging, which is much more flexible and can learn nonlinearities and interactions between explanatory variables from the data. More specifically, the MLP is trained to learn the function $g()$, which takes the model probabilities, $P_{n,m}$ and explanatory variables $Z_{n}$ to produce the model weight: ${\pi}_{i,n,m}=g(\omega\mid P_{n,m},z_{n})$.
3 Case studies
In this paper, we illustrate the performance of the proposed approach in two case studies. In what follows, we first describe the two datasets, before talking about how the data was divided for the purpose of the present paper. We then talk about the individual sub-models and the specification and estimation of the model averaging structure. We finally look at the results.
3.1 Data
We use two different revealed preference datasets in this study, both focussing on mode choice.
3.1.1 DECISIONS data
The first dataset comes from a large-scale survey conducted as part of the DECISIONS project carried out by the Choice Modelling Centre at the University of Leeds (Calastri et al., 2020). We specifically use the GPS tracking component of this survey, where the data used for this work corresponds to the observed mode choice behaviour.
After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 12,524 trips made by 540 individuals remained. For each trip, individuals travelled by one of six modes: car (48%), bus (15%), rail (5%), taxi (3%), cycling (3%) or walking (26%). The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
The attributes of the alternatives used in the models for the present paper include in-vehicle travel time, out-of-vehicle travel time, and travel cost.
3.1.2 LPMC data
The second dataset is the London mode choice data compiled by Hillel et al. (2018), referred to hereafter as LPMC (London passenger mode choice).
We use a sample of $80,943$ trips made by $31,921$ individuals. This dataset contains four alternatives: walking (18%), cycling (3%), public transport (35%, grouping together bus and rail) and driving (44%, grouping together private car and taxi rides). The data again covers a very wide range of trip distances, going from just under 100 metres to just over $40$ kilometres, with a mean distance of $4.6$ kilometres and a median distance of $2.8$ kilometres.
Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two socio-demographic variables, namely car ownership and driving license status.
3.2 Estimation and validation data
As outlined in the earlier discussions, our approach to model averaging relies on the idea of excluding part of the distance distribution in the training of the individual sub-models, by setting $d_{a}>d_{min}$ and $d_{b}<d_{max}$. For the purpose of testing the performance of the model averaging approach, we go one step further in our empirical work, by excluding very short and very long trips from the training of the model averaging structure as well, thus allowing us to validate the out-of-distribution performance.
Formally, we have that the range of distances in the data is given by $D_{data}=\left[d_{min},d_{max}\right]$. The range used in training the individual sub-models is given by $D_{train,sub}=\left[d_{a,sub},d_{b,sub}\right]$. Finally, the model averaging models are estimated on $D_{train,MA}=\left[d_{a,MA},d_{b,MA}\right]$. Overall, the relationship holds that $d_{min}<d_{a,MA}<d_{a,sub}<d_{b,sub}<d_{b,MA}<d_{max}$. For the purposes of our analysis, we set $d_{a,MA}$ and $d_{b,MA}$ to be at the 10th and 90th percentiles of the sample distance distribution $D_{data}$, while $d_{a,sub}$ and $d_{b,sub}$ are set at the 20th and 80th percentiles of the sample distance distribution $D_{data}$.
In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample validation [^1] The splitting of the sample into estimation, validation (and additionally test data for MLP and XGB) was done at the observation rather than person/household level. This might be seen as a limitation, however, none of the models we used explicitly accommodated any individual-specific effects, and information on IDs was not provided to the MLP and XGB training. We thus believe the impact of the way we have split the data to be negligible.. The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1, where the larger sample validation size for the two extreme segments is a result of this data being used for validation only. In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work and would likely also use a smaller validation sample - we made these decisions with a view to validating the model idea.
Figure 1: Estimation and validation data
Table 1: Estimation and validation data sample sizes for case studies
DECISIONS data LPMC data
Distance (km) Estimation Validation Distance (km) Estimation Validation
Individual MA Individual MA
out-of-distribution for MA (below, i.e. $d_{min}tod_{a,MA}$) <1.33 0 0 1,253 <0.75 0 0 8,092
out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 1,026 228 0.75 to 1.11 0 6,508 1,580
Within distribution for MA and sub-models 1.95 to 11.95 5,952 1,560 1.11 to 7.41 38,855 9,696
out-of-distribution for sub-models (above, i.e. $d_{b,sub}$ to $d_{b,MA}$) 11.95 to 19.52 0 993 259 7.41 to 11.22 0 6,481 1,623
out-of-distribution for MA (above, i.e. $d_{b,MA}$ to $d_{max}$) 19.52 to 105.9 0 0 1,253 11.22 to 40.1 0 0 8,108
Total 5,952 2,019 4,553 38,855 12,989 29,099
3.3 Model types
The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning. Many other model types could have been considered [^2] A notable omission in our list are mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out. The only exception comes with the use of posterior distributions which are not applicable in the context of out-of-sample prediction. - we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
Nested Logit (NL):
Our NL models (cf. Daly and Zachary, 1978) use a full set (i.e. $J-1$) of alternative specific constants (ASC). For continuous attributes, we use a linear plus logarithmic specification, i.e. $V=...+\beta_{x}x+\beta_{log-x}log(x)$ to capture non-linearity in sensitivities. For the DECISIONS data, we use three nests, grouping car together with taxi, bus with train, and walking with cycling. For the LPMC data, we use two nests, grouping car together with bus, and walking together with cycling, but this structure ended up collapsing to a Multinomial Logit (MNL) model.
Decision field theory (DFT):
DFT is a dynamic, stochastic model, introduced by Busemeyer and Townsend (1993). The key idea of the DFT model is that the preferences for different alternatives update over time whilst the decision-maker considers the different alternatives and their attributes. We use the implementation of Hancock et al. (2021), with the valence functions following the same specifications as the utility in NL.
MultiLayer Perceptron (MLP):
This model comprises an input layer with input nodes, one or more hidden layers with hidden nodes, and an output layer with output nodes. In this model, signals propagate forward through the links connecting the nodes. The links have numeric weights $w$, which are learned from the data. At each link, the weights are multiplied by the input value from the previous nodes. At the node, the sum of the inputs is taken, and an activation function is applied. The outcome is propagated to the nodes in the next layer. We use tanh activation functions in all hidden layers, except the final one where a linear layer is used. In the output layer, a Softmax function (i.e. a logit) is applied to produce choice probabilities for each alternative.
XGBoost (XGB):
The XGB model comprises a series of sequentially applied decision trees. A decision tree is a sequence of simple IF-THEN rules, optimised to classify data accurately. In the XGB model, each decision tree in the series ‘corrects’ the mispredictions of the models before it. This process is referred to as ‘boosting’. The term ‘gradient’ refers to the fact that each successive tree is fitted using the gradient of the loss function—effectively via gradient descent. Unlike standard Gradient Boosting Machines (GBM), XGBoost includes further optimisations, such as second-order gradient information and advanced regularisation, to improve speed and mitigate overfitting.
The NL and DFT models were coded and estimated using Apollo (cf. Hess and Palma, 2019), while the MLP and XGB models were implemented in Python using the SciKit-learn (Pedregosa et al., 2011) and XGBoost (Chen and Guestrin, 2016) libraries respectively. A hyperparameter search was conducted for both models to optimise their performance.
For the MLP, the hyperparameter search considered the number of hidden layers, the number of nodes per layer, the learning rate, and the L2 regularisation strength. The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the learning rate controls how quickly the model updates its weights during training. L2 regularisation helps prevent overfitting by penalising large weight values. Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration. The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, and an L2 regularisation strength of 0.1. Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged their predictions to account for this variability.
For the XGBoost model, the hyperparameter search focused on selecting a subset of features for each tree split, determining the minimum loss reduction needed for further node splitting, setting the maximum tree depth, and specifying the minimum sum of instance weights required in a child node. The first controls the diversity of the model by randomly selecting a fraction of features at each split. The second acts as a pruning mechanism by restricting splits that do not provide sufficient improvement. The third helps balance model complexity and the risk of overfitting. The fourth serves as a regularisation measure, preventing excessively complex splits by enforcing a minimum requirement for node creation. The final parameter determines the fraction of training samples used in each boosting iteration. Similar to the MLPs, since the optimal hyperparameters were nearly identical for both datasets, we opted to use a single configuration. The best-performing values selected were a feature sampling rate of 0.7, a pruning threshold of 2, a maximum tree depth of 2, a minimum child node weight requirement of 1, and a subsampling rate of 1333This is rather shallow, especially in comparison with Hillel (2021), but was the outcome of the hyperperformance tunning on the subset of the data we used.. As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training. This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
3.4 Model averaging: specification and estimation
Unlike the MLP models implemented for the training on the data sets (see Section 3.3), the MLP for the model averaging is implemented in PyTorch Paszke et al. (2019). This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in SciKit-learn. To find the best hyperparameters, hyperparameter tuning was conducted using the following parameters: batch size, number of hidden layers and nodes, learning rate and L2 regularisation. For the DECISIONS data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.0001. For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
We repeated the training 100 times using the optimal configuration. As we aim for high prediction performance, we retained only the top 20 percent of models [^4] This selection was based on the log-likelihood. For this, we used the full data, i.e. training and test data, but excluded the validation data., discarding those that became trapped in poor local maxima. This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
Figure 2 shows the structure for our model averaging approach. As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model. The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model averaging MLP.
Figure 2: Model averaging structure
3.5 Empirical results
In this section, we present the empirical results for the two case studies. We first focus on the performance of the individual sub-models, comparing this across distance segments and also discussing out-of-sample performance. We then look at the results for the model averaging structure.
3.5.1 Performance of individual sub-models
Table 2 presents the performance of the different models on the estimation sample, covering distance segments (deciles) 3 to 8, as illustrated in Figure 1.
We see that for the DECISIONS data, DFT obtains a higher log-likelihood (LL) than NL, while MNL outperforms DFT in the LPMC data. In both studies, the two machine learning approaches obtain much higher LL than NL/MNL and DFT. Proportionally, the difference is larger in the DECISIONS case study than in the LPMC one, where in the former, MLP also has a bigger advantage over XGB. No formal statistical tests are carried out to compare the different model structures as these are not applicable to the machine learning structures.
Table 2: Final log-likelihood of different model structures (covering distance segments 3-8) of the estimation set
Model DECISIONS LPMC
NL/MNL -2,540.83 -28,782.34
DFT -2,522.93 -28,794.57
MLP -2,095.93 -27,605.21
XGB -2,254.93 -27,635.04
As a next step, Figure 3 looks at the performance of these models across the different distance segments obtained by splitting the data into deciles of the distance distribution. We do this separately for the estimation and validation data. A point to note when studying these results is that for the individual models, only segments 3 to 8 were used in estimating the model parameters, but the probabilities in segments 2 to 9 for the individual models also contributed to the model averaging, as explained in Figure 1. The values shown for segments 2 and 9 in the left two graphs in Figure 3 thus in effect also relate to out-of-distribution validation for the individual models.
Figure 3: Estimation and validation performance
Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 2 apply in each of the 6 distance segments used in estimation (i.e. segments 3-8). Across the distance segments, there are some where XGB outperforms MLP (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases. These findings already show that different models perform differently well across trip distances. Interestingly, while the LPMC results show that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
We next turn to segments 2 and 9 in the left two graphs, giving us initial insights into out-of-distribution performance, remembering that the values shown here are the predicted probabilities for the chosen alternative, using the model parameters obtained when estimating the models on segments 3 to 8. In the DECISIONS data, we see all models perform very similarly in segment 2 (except for XGB), while in segment 9, NL clearly outperforms the data-driven models (MLP and XGB) and DFT. In the LPMC models, we see that MLP performs best in both segments 2 and 9, but the differences are very small.
We next turn to the validation results, shown in the right two graphs in Figure 3. We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data. For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for the NL and DFT, with an average gain in per observation LL of $0.02752$ and $0.01811$, respectively, where this is most notable in segment 8. For MLP and XGB, we see small drops, with changes of $-0.0134$ and $-0.0058$, respectively. In the LPMC models, the changes are smaller, with gains in per observation LL of $0.0098$ for MNL, $0.01$ for DFT, $0.0043$ for MLP, and a drop by $-0.0008$ for XGB. Overall, these results raise no concerns about overfitting to the estimation data, at least within the distribution of distances covered by the estimation data.
The findings for segments 2 and 9 on the validation data largely confirm those we saw in the two graphs on the left in Figure 3. We finally turn to segments 1 and 10. From a behavioural perspective, we see that for the DECISIONS data, mode choice on very short tips is much easier to predict. For short trips in the LPMC dataset, there is a clear advantage for the behavioural models. For the longest trips (segment 10), the trend is less clear. In the DECISIONS data, MLP performs comparatively well. In the LPMC data, both data-driven models do comparatively well, while MNL suffers a big drop.
3.5.2 Model averaging results
In model averaging, data covering segments 2-9 were used, but with model parameters for the individual models estimated on segments 3-8 only (cf. discussion around Figure 1). Table 3 shows the performance (in terms of LL on the estimation set) of the four individual models, alongside that of the model averaging (MA) structure. In both datasets, the MA structure clearly obtains a higher LL than any of the four individual models, although in LPMC, the difference to the best-fitting sub-model (MLP) is only around 0.13%. This in itself is simply a reflection of the overall dominance of MLP over the other three model types in segments 2-9 for this data, and further insights will be gained in model validation. A potential reason for this dominance could be the larger sample size compared to the DECISIONS data, which could enable MLP and XGB to better learn the data generating process.
Table 3: Log-Likelihood for model averaging: individual components and overall structure (covering distance segments 2-9), with gains by MA over individual models shown in brackets
Model DECISIONS LPMC
NL/MNL -3,397.316 (13.62%) -38,383.5 (3.91%)
DFT -3,451.558 (14.98%) -38,366.13 (3.86%)
MLP -3,019.805 (2.83%) -36,931.39 (0.13%)
XGB -3,202.91 (8.38%) -37,228.17 (0.92%)
MA -2,934.45 -36,884.39
The inclusion of segments 2 and 9 allows the model averaging structure to learn about out-of-distribution performance and hence guide the process of linking the weights given to individual models to the difference between the distance for that trip and the data used in estimation (i.e. segments 3 to 8). As explained in Section 3.4, a very flexible neural network was used to estimate this relationship, with no a priori assumptions about the functional form. The resulting relationship is illustrated in Figure 4.
For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e. distances on which the individual sub-models were estimated. In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP. These findings are entirely in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
Once we go outside the $D_{train,sub}$ interval, the pattern changes. In both datasets, we observe that for trips with distances below $d_{a,sub}$ (i.e. segments 1&2), the weight for behavioural models increases. In the DECISIONS data, only a small share goes to NL, with quite equal weight for the other three models. On the other hand, in LPMC, MNL essentially obtains all the weight. For longer trips, i.e. above $d_{b,sub}$, NL gets the vast majority of the weight in the DECISIONS data, with the rest going to DFT. For LPMC, MLP retains a large weight, but this decreases while the weight for MNL and DFT increases with distance.
Figure 4: Model averaging: weights given to individual models
The final step in our analysis of results concerns the performance of the model averaging structure on the validation data. Table 4 shows the LL on the validation data, covering the entire range of distances, for the four sub-models and for model averaging. As in estimation, we see that the model averaging structure obtains the least negative LL. What is striking however is the contrast with the results on the estimation data, i.e. Table 2. With the exception of DFT on the LPMC data, the gains by model averaging over the sub-models are larger on the validation data than on the estimation data. This provides initial evidence to support the notion that model averaging is less susceptible to over-fitting on the estimation data than data-driven models.
Table 4: Log-Likelihood for model averaging: individual components and overall structure on validation data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
Model DECISIONS LPMC
NL/MNL -1,948.29 (15.32%) -22,233.56 (14.1%)
DFT -1,980.411 (16.7%) -19,414.46 (1.62%)
MLP -1,722.19 (4.21%) -19,869.84 (3.88%)
XGB -1,895.97 (12.99%) -22,131.08 (13.7%)
MA -1,649.75 -19,099.30
There are two possible influences that could drive the differences between Table 3 and Table 4. The first of these is the potential that model averaging reduces the risk of over-fitting to the estimation data, which can be tested by comparing the performance on segments 2-9 between the estimation and validation data. The second is the potential that model averaging offers improved out-of-distribution performance, which can be tested by comparing the performance on segments 1 and 10 between the sub-models and the model averaging structure.
To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and validation. We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 7 out of 8 for the LPMC data. When we turn to the validation data, model averaging performs best (or joint best) in 9 out of 10 segments for the DECISIONS data, and 8 out of 10 segments for the LPMC data. These per segment statistics go hand in and hand with the best overall LL obtained by model averaging as reported in Table 3 and Table 4.
Figure 5: MA Estimation and validation performance
In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments where the per observation LL increases and segments where it decreases. No clear pattern emerges in the comparison across models and segments. Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.3419 vs -0.3628 for MLP). For the LPMC data, a similar story emerges. Model averaging is outperformed slightly by NL in segment 1, and slightly by DFT in segment 10, but overall performs the best of all models when looking at segments 1 and 10 together (-0.6078 vs -0.6153 for DFT).
Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance (when looking at segments 2-9), aside from the potential suggestion that the risk of overfitting is reduced with more behavioural models, as already shown in Figure 3. There is no particular benefit for model averaging over other models in this context (given the dominance of the data-driven models), and nor was this the expectation. On the other hand, there is clear evidence that our model averaging approach offers benefits in terms of out-of-distribution performance, which was its intended aim.
4 Conclusions
This work has taken an important step forward in combining insights from different modelling approaches for travel demand forecasting. Specifically, we have shown that different models predict choices differently well depending on how far away from the estimation sample the prediction takes place. This result is not surprising in itself but is quantified by our work.
The main contribution comes in the development of an MLP-based model averaging approach that estimates weights for different models as a function of the distance away from the estimation data. Across two different case studies looking at mode choice, we see that data-driven machine learning approaches obtain almost the entire weight when making predictions for trips in areas similar to the trips in the estimation data. However, once we move outside that area and look at out-of-distribution prediction, the model averaging approach increases the weight to the models with behavioural underpinnings, sometimes substantially so.
In terms of prediction performance, we observe that model averaging offers the best overall performance on both the estimation and validation data. Crucially, this advantage is at its largest when making out-of-distribution predictions, in our case, predicting mode choice for trips whose distance is outside the range covered by the estimation data.
There are, of course, some opportunities for further refinement.
First, in our implementation, we used the same data to estimate the sub-models and the model averaging. In case of overfitting by one of the sub-models (an issue that did not arise in our case), this could create the risk that model averaging assigns disproportional weight to that model, which would jeopardise the benefits of the approach. In studies with larger sample sizes, a further splitting of the data could be considered.
Second, in applying our proposed model averaging approach, an analyst needs to define what distribution means. In our work, we have focussed on the obvious example of trip distance. A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data. Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time.
Finally, we used MLP for the model averaging structure. Other meta-models, such as hybrid architectures approaches like TasteNetMNL (cf. Han et al., 2022), are an interesting avenue for future research.
Acknowledgements
Stephane Hess acknowledges the support of the Delft Excellence Fund, as well as the European Research Council through the advanced grant 101020940-SYNERGY. Sander van Cranenburgh acknowledges the support of the TU Delft AI Labs programme. The authors are thankful to Thomas Hancock, Tim Hillel and Georges Sfeir for comments on an earlier draft.